Suppose a simple random sample of size n=50 is taken from a population of N=15,000 with a population proportion of p=0.4 having a certain characteristic. What is the probability that a simple random sample will have x > 21 having that characteristic?

(Round to four decimal places)
1
Expert's answer
2020-08-03T18:43:20-0400

Check that the sample size is large enough:


"n\cdot p=50\cdot0.4=20\geq10"

"np(1-p)=50\cdot0.4(1-0.4)=12\geq10"

Sampling distribution for "\hat{p}" is approximately normal:


"\mu_{\hat{p}}=p=0.4"

"\sigma_{\hat{p}}=\sqrt{\dfrac{p(1-p)}{n}}=\sqrt{\dfrac{0.4(1-0.4)}{50}}\approx0.069282"

"{21\over 50}=0.42"

"P(\hat{p}>0.42)=1-P(\hat{p}\leq0.42)="

"=1-P(Z\leq\dfrac{0.42-0.4}{0.069282})\approx1-P(Z\leq0.288675)\approx"

"\approx0.386415"

The probability that a simple random sample will have x > 21 having that characteristic is approximately "0.3864."